# Interactive Inference: A Neuromorphic Theory of Human-Computer Interaction

### v4

@@ Discussion
The above results are a first step towards empirical validation of both our capacity and error models. Results show a good to excellent fit between between predicted and observed models of the task. Driver information processing capacity during deceleration events in a simulated car following task can be measured in real time using the simplified KL divergence of the noise-to-signal ratio of the distance to the car followed, over a time window. Our hypotheses that both distance $S$ and its variance NN to the lead car significantly affect the amount of information consumed by the driver during such events were confirmed, with more information consumed at lower $S$s and higher NNs.
@@
Qualitative ratings by participants were in line with the above. They rated the task as more difficult with higher noise-to-signal ratios (and thus IDIDs), albeit with bipolar ratings of middle IDIDs. We attribute this to participants perceiving distant targets (S>4.8m)(S>4.8~m) as easier to process than close targets with similar signal-to-noise ratios, perhaps due to the 2D nature of our simulator setup. We again note that our results are limited to modeling driver responses to lead car decelerations in a simulated car following task using specific road conditions, car physics, and a sample set of participants.
@@ Conclusions
Neuromorphic Human-Computer Interaction aims to improve user interfaces by exploring how concepts from neuroscience might inform design. Interactive Inference offers a simplified interpretation of the neuroscience theory of Active Inference in user interface design. It postulates that users perform Bayesian inference on current progress and goal distributions to learn tasks and predict their next action. The Bayesian surprise (relative entropy) between goal and progress distributions can be modeled as a square function. However, human capacity to process Bayesian surprise follows a logarithmic function. This explains why errors occur when the user runs out of capacity. We believe our model may allow the quantitative analysis of performance and error in a single framework that allows estimation of the mental load of a task and showed how three basic laws of HCI, Hick’s Law, Fitts’ Law and the Power Law of Practice can be derived from Interactive Inference. We evaluated the use of the model in an empirical study. In our simulated car following task, driver information processing capacity during braking events was linearly related, with good fit, to a logarithmic model of the noise-to-signal ratio of the distance to a lead car, plus 1 (in bits). We found a significant effect of both distance $S$ and variation NN on capacity, with a mean driver Index of Performance of almost 4 bits of difficulty (as defined by $log_{2}(N/S+1))$ per bit of information. Participants’ qualitative ratings followed the same trend, with significantly higher ratings of difficulty with higher $N/S$ ratios. We used our model to derive a remaining capacity for the driver, by subtracting the log-linear curve from mean capacity, in bits. We also fit error (collision) rates, with excellent fit, to the square of the signal-to-noise ratio of the distance to the lead car, in bits. We hypothesize that collisions are due to unprocessed information where driver capacity is exceeded. Overall, the empirical results provide support for our Interactive Interference model of HCI. While car-following is just one case study, it provides a framework for conducting further empirical research in Interactive Inference, and evidence that such research is well worth doing. We invite other researchers to test and apply this model to other tasks, to see how well these models predict human performance, and to confirm or refute the general applicability of Neuromorphic Human-Computer Interaction in general, and Interactive Inference in particular.

### v5

@@ Discussion
The above results are a first step towards empirical validation of both our capacity and error models. Results show a good to excellent fit between between predicted and observed models of the task. Driver information processing capacity during deceleration events in a simulated car following task can be measured in real time using the simplified KL divergence of the noise-to-signal ratio of the distance to the car followed, over a time window. Our hypotheses that both distance $S$ and its variance $N$ to the lead car significantly affect the amount of information consumed by the driver during such events were confirmed, with more information consumed at lower $S$s and higher $N$s.
@@
Qualitative ratings by participants were in line with the above. They rated the task as more difficult with higher noise-to-signal ratios (and thus $ID$s), albeit with bipolar ratings of middle $ID$s. We attribute this to participants perceiving distant targets (S>4.8m)(S>4.8~m) as easier to process than close targets with similar signal-to-noise ratios, perhaps due to the 2D nature of our simulator setup. We again note that our results are limited to modeling driver responses to lead car decelerations in a simulated car following task using specific road conditions, car physics, and a sample set of participants.
@@ Conclusions
Neuromorphic Human-Computer Interaction aims to improve user interfaces by exploring how concepts from neuroscience might inform design. Interactive Inference offers a simplified interpretation of the neuroscience theory of Active Inference in user interface design. It postulates that users perform Bayesian inference on current progress and goal distributions to learn tasks and predict their next action. The Bayesian surprise (relative entropy) between goal and progress distributions can be modeled as a square function. However, human capacity to process Bayesian surprise follows a logarithmic function. This explains why errors occur when the user runs out of capacity. We believe our model may allow the quantitative analysis of performance and error in a single framework that allows estimation of the mental load of a task and showed how three basic laws of HCI, Hick’s Law, Fitts’ Law and the Power Law of Practice can be derived from Interactive Inference. We evaluated the use of the model in an empirical study. In our simulated car following task, driver information processing capacity during braking events was linearly related, with good fit, to a logarithmic model of the noise-to-signal ratio of the distance to a lead car, plus 1 (in bits). We found a significant effect of both distance $S$ and variation $N$ on capacity, with a mean driver Index of Performance of almost 4 bits of difficulty (as defined by $log_{2}(N/S+1))$ per bit of information. Participants’ qualitative ratings followed the same trend, with significantly higher ratings of difficulty with higher $N/S$ ratios. We used our model to derive a remaining capacity for the driver, by subtracting the log-linear curve from mean capacity, in bits. We also fit error (collision) rates, with excellent fit, to the square of the signal-to-noise ratio of the distance to the lead car, in bits. We hypothesize that collisions are due to unprocessed information where driver capacity is exceeded. Overall, the empirical results provide support for our Interactive Interference model of HCI. While car-following is just one case study, it provides a framework for conducting further empirical research in Interactive Inference, and evidence that such research is well worth doing. We invite other researchers to test and apply this model to other tasks, to see how well these models predict human performance, and to confirm or refute the general applicability of Neuromorphic Human-Computer Interaction in general, and Interactive Inference in particular.